You believe the stock in Freeze Frame Co. is going to fall, so you short 950 shares at a price of $74. The initial margin is 50 percent.

1.

Construct the equity balance sheet for the original trade. (Input all amounts as positive values. Omit the "$" sign in your response.)

2-a.

Construct an equity balance sheet for a stock price of $65 per share. (Input all amounts as positive values. Omit the "$" sign in your response.)

What is your margin? (Round your answer to 2 decimal places. Omit the "%" sign in your response.)

What is your effective annual return if you cover your short position at this price in 6 months? (Round your answer to 2 decimal places. Omit the "%" sign in your response.)

Construct an equity balance sheet for a stock price of $79 per share.(Input all amounts as positive values. Omit the "$" sign in your response.)

What is your margin? (Round your answer to 2 decimal places. Omit the "%" sign in your response.)

What is your effective annual return if you cover your short position at this price in 6 months? (Negative amounts should be indicated by a minus sign. Round your answer to 2 decimal places. Omit the "%" sign in your response.)
Answer #1
1
Equity Balance sheet
Assets Liabilities and account equity
Cash from sale (950*74) 70300 Short position 70300
Initial margin deposit (70,300*50%) 35150 Account equity 35150
Total 105450 105450
2-a.
Equity Balance sheet
Assets Liabilities and account equity
Cash from sale (950*74) 70300 Short position (950*65) 61750
Initial margin deposit (70,300*50%) 35150 Account equity (105450-61750) 43700
Total 105450 105450
With change in stock price of share the short position and account equity would change
Calculation of margin
Margin % = Account equity/Short position
Margin % = 43,700/61,750
Margin % 70.77%
The margin is 70.77
The 6 month return = (43,700-35,150)/35,150 24.32%
Effective annual return = (1+0.2432)^(12/6) - 1 54.55%
The effective annual return if short position is covered at this price in 6 months is 54.55
Construct an equity balance sheet for a stock price of $79 per share
Equity Balance sheet
Assets Liabilities and account equity
Cash from sale (950*74) 70300 Short position (950*79) 75050
Initial margin deposit (70,300*50%) 35150 Account equity (105450-75050) 30400
Total 105450 105450
Calculation of margin
Margin % = Account equity/Short position
Margin % = 30,400/75,050
Margin % 40.51%
The margin is 40.51
The 6 month return = (30,400-35,150)/35,150 -13.51%
Effective annual return = (1-0.1351)^(12/6) - 1 -25.19%
The effective annual return if short position is covered at this price in 6 months is -25.19